# Picky elements, subnormalisers, and character correspondences

### v1

@@ 5.1. Semisimple picky elements
The centraliser ${\mathbf{C}}:={\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is an $F$-stable Levi subgroup of ${\mathbf{G}}$. Its structure is described in [14, Exmp. 3.5.14 and 3.5.15] for ${\mathbf{G}}$ of classical type. Note that since ${\mathbf{S}}$ is a Sylow $e$-torus, it has the form ${\mathbf{T}}{\mathbf{H}}$ with ${\mathbf{T}}$ a torus whose order polynomial only involves factors $\Phi_{d}$ with $d\leq 2e$, and ${\mathbf{H}}$ a semisimple group of rank less than $e$. The claim can easily be checked by inspection of the order formulas (e.g. in [24, Tab. 24.1]). For groups of exceptional type, Table 3.3 in [14] shows that ${\mathbf{C}}$ is itself a torus, with the required property, unless either $e=4$ and ${\mathbf{G}}$ has type $E_{7}$, but the claim still holds in that case, or ${\mathbf{G}}^{F}={}^{3}\!D_{4}(q)$ and $\ell=3$ (where the claim fails). ∎
@@
Assume ${\mathbf{G}}$ is simple. Assume the order of the relative Weyl group $W_{e}$ of a Sylow $e$-torus of $G$ is divisible by $\ell$, where $\ell>2$. Then the normaliser of a Sylow $\ell$-subgroup of $W_{e}$ contains no elements of order $\ell r$ for primes $r\equiv 1\pmod{\ell}$.
@@
[…]
First assume $x\in{\mathbf{S}}_{\ell}^{F}$. By Proposition 5.5, since $\ell\geq 5$, $W_{e}$ has more than one Sylow $\ell$-subgroup, so $x$ lies in two different Sylow $\ell$-subgroups of ${\mathbf{N}}_{G}({\mathbf{S}})$ and hence of $G$. Now assume $x$ does not lie in any Sylow $e$-torus, and let ${\mathbf{T}}$ be an $F$-stable torus of ${\mathbf{G}}$ containing $x$. Then the order polynomial of ${\mathbf{T}}$ must be divisible by a cyclotomic polynomial $\Phi_{e\ell^{i}}$ for some $i\geq 1$ and hence $|{\mathbf{C}}_{G}(x)|$ is divisible by $\Phi_{e\ell^{i}}(q)$. Let $r$ be a Zsigmondy primitive prime divisor of $\Phi_{e\ell^{i}}(q)$, which exists as $e\ell^{i}$ is divisible by a prime at least 5. On the other hand, by Proposition 5.4, the order polynomial of ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is not divisible by $\Phi_{e\ell^{i}}$, so $|{\mathbf{C}}_{G}({\mathbf{S}})|$ is prime to $r$. Thus $W_{e}$ must contain an element of order $\ell r$. Now note that $r\equiv 1\pmod{\ell}$. Thus Lemma 5.6 together with Lemma 2.1 show that $x$ cannot be picky in ${\mathbf{N}}_{G}({\mathbf{S}})$ and so neither in $G$. ∎
@@
[…]
@@
[…]
@@
Next, the exceptions of Proposition 5.5 occur precisely for the groups ${\operatorname{SL}}_{3}(\epsilon q)$, $G_{2}(q)$ and ${}^{3}\!D_{4}(q)$, since these are the only groups with a relative Weyl group of a Sylow $e$-torus isomorphic to ${\mathfrak{S}}_{3}$ or $W(G_{2})$. If $G={\operatorname{SL}}_{3}(\epsilon q)$ we have $3|(q-\epsilon)$ since we assume Sylow 3-subgroups of $G$ are non-abelian, and even $9|(q-\epsilon)$ by the previous paragraph. Then, by [23, Thm 5.16] the normaliser of a Sylow 3-subgroup $P$ of $G$ lies inside the normaliser ${\mathbf{N}}_{G}({\mathbf{S}})$ of a Sylow $e$-torus ${\mathbf{S}}$. If $x$ is a 3-element not lying in a conjugate of ${\mathbf{S}}_{3}^{F}$, then we may conclude as in the proof of Theorem 5.7 that $x$ is not picky. If $x\in{\mathbf{S}}_{3}^{F}$ is not regular, its centraliser involves an $A_{1}$-type group, hence is not contained in ${\mathbf{N}}_{G}({\mathbf{S}})={\mathbf{S}}^{F}.{\mathfrak{S}}_{3}$ and again $x$ is not picky by Lemma 2.1. Now assume $x\in{\mathbf{S}}_{3}^{F}$ is regular. If ${\mathbf{N}}_{G}({\mathbf{S}})$ has more than one Sylow 3-subgroup, again $x$ is not picky. If ${\mathbf{N}}_{G}({\mathbf{S}})$ has a unique Sylow 3-subgroup, that is, the Sylow 3-subgroup is normal, then the elements of order three in $W_{e}$ need to centralise the $3^{\prime}$-part of ${\mathbf{S}}^{F}$ which by inspection forces ${\mathbf{S}}^{F}$ to be a 3-group. Thus $q-\epsilon$ is a power of 3 which together with $q\equiv\epsilon\pmod{9}$ forces $q=8$, $\epsilon=-1$. In the latter case, all regular $x\in{\mathbf{S}}_{3}^{F}$ are picky by direct computation, as claimed in (1).
@@
Next, if $G={}^{3}\!D_{4}(q)$ with $q\equiv\epsilon\pmod{3}$, $\epsilon\in\{\pm 1\}$, then the normaliser of a Sylow 3-subgroup is contained in a torus normaliser of the form $N=(q^{3}-\epsilon)(q-\epsilon).W(G_{2})$ (see the discussion in the proof of [23, Thm 5.14]). Now elements of order 3 in $W(G_{2})$ act like a field automorphism on the torus of order $q^{3}-\epsilon$. In particular, as soon as $q^{2}+\epsilon q+1$ has a primitive prime divisor (necessarily distinct from 3), a Sylow 3-subgroup is not normal in $N$ and we may conclude as before. This leaves $q=2$, so $G={}^{3}\!D_{4}(2)$. Here, by explicit computation, the elements of order 9 with centraliser $A_{1}(q).(q^{3}+1)$ are picky.
Finally, if there exists no primitive prime divisor for $\Phi_{e3^{i}}(q)$ with $i\geq 1$, then $i=1$, $e=q=2$. Now by inspection, the order polynomial of the centraliser of a Sylow 2-torus for a simple group ${\mathbf{G}}$ is never divisible by $\Phi_{3}$ (again using [14, Exmp. 3.5.14, 3.5.15, and Tab. 3.3]) and hence the argument used in the generic case applies here with the prime divisor $r=7$ of $\Phi_{6}(2)$ to show there do not arise picky elements from this situation. ∎
Observe that for the first group in Case (1) of Theorem 5.8 the Sylow 3-subgroups of the ambient simple group $G/{\mathbf{Z}}(G)$ are abelian and $x$ is regular, while in Case (3), the simple group $[G,G]$ is a Lie type group in characteristic 3.
The case $\ell=2$ is considerably more involved. We have not (yet) been able to obtain a definite answer and the results will by all appearance be less clean than for odd $\ell$, involving Fermat and Mersenne primes, for example. We therefore postpone it to a later investigation.
5.2. The Suzuki and Ree groups
@@
We now assume that ${\mathbf{G}}$ is of type $B_{2}$, $G_{2}$ or $F_{4}$ in characteristic $p=2,3$, or 2, respectively, and $F$ is a Steinberg endomorphism of ${\mathbf{G}}$ such that $G={\mathbf{G}}^{F}$ is a Suzuki or Ree group. As before, $\ell$ denotes a prime distinct from the defining characteristic $p$ of ${\mathbf{G}}$. We obtain the analogue of Theorems 5.2 and 5.7:
Proposition 5.9.
@@
In the above setting, an $\ell$-element $x\in G$ is picky if and only if $\ell>3$ and $x$ is regular.
@@
First assume that $\ell>3$. Then $\ell$ does not divide the order of the Weyl group of ${\mathbf{G}}$ and hence the Sylow $\ell$-subgroups of $G$ are abelian (see [2, Cor. 3.13, p.259]). We may now proceed as in the proof of Theorem 5.2, replacing cyclotomic polynomials over ${\mathbb{Q}}$ by suitable cyclotomic polynomials over ${\mathbb{Q}}(\sqrt{p})$, with corresponding Sylow tori, for which the Sylow theorems continue to hold, see [14, 3.5.3, 3.5.4].
[…]
Finally, $\ell=2$ is non-defining only for $G={}^{2}G_{2}(q^{2})$. Here, Sylow 2-subgroups are elementary abelian of order 8 and there is a unique class of involutions, with centraliser ${\operatorname{PSL}}_{2}(q^{2})\times 2$. Since the Sylow 2-normaliser, of structure $7.32^{3}.$7.3, does not contain such a subgroup, the involutions cannot be picky by Lemma 2.1. ∎
5.3. On the conjecture for non-defining primes
Lemma 5.10.
Conjecture 1+ holds for the picky $3$-elements of ${\operatorname{SU}}_{3}(8)$, $G_{2}(8)$, ${}^{3}\!D_{4}(2)$ and of $G_{2}(2)$.
Proof.
Let $G$ be one of the four groups in question, $P\leq G$ a Sylow 3-subgroup and $x\in P$ picky. By direct computation in GAP, for ${\operatorname{PSU}}_{3}(8)$ there are six irreducible $3^{\prime}$-characters and three characters with $\chi(1)_{3}=3$ for $G$ as well as for ${\mathbf{N}}_{G}(P)$ that do not vanish on $x$, with matching character fields and 3-parts of values. For ${}^{3}\!D_{4}(2)$ there are nine $3^{\prime}$-characters and six characters with $\chi(1)_{3}=3$ for $G$ as well as for ${\mathbf{N}}_{G}(P)$ that do not vanish on $x$. For $G_{2}(2)$ there are nine $3^{\prime}$-characters and five characters with $\chi(1)_{3}=3$ for $G$ as well as for ${\mathbf{N}}_{G}(P)$ that do not vanish on the 3-central 3-elements; only the first nine do not vanish on the regular 3-elements of $G^{\prime}={\operatorname{SU}}_{3}(3)$. Again Conjecture 1+ is satisfied.
[…]

### v2

@@ 5.1. The case of abelian Sylow $\ell$-subgroups
The precise determination of subnormalisers in the abelian Sylow case seems to require a classification of the overgroups of normalisers of Sylow $d$-tori for simple algebraic groups, a non-trivial and interesting problem, as the following example shows (see also Corollary 6.9 in the algebraic group case):
Example 5.5.
Let $\mathbf{G}$} be of type $F_{4}$ and $\ell>3$ dividing $\Phi_{3}(q)$. Let $x\in G={\mathbf{G}}^{F}$ be an $\ell$-element with centraliser $A_{2}(q).\Phi_{3}$. There is a subgroup $^{3}\!D_{4}(q).3$ of $G$ containing the normalisers of all Sylow $d$-tori of ${\mathbf{C}}_{G}(x)$, and hence also ${\operatorname{Sub}}_{G}(x)$ by Proposition 5.2. But it is not immediately obvious that the subnormaliser could not be smaller.
5.2. Picky elements in the non-abelian case
@@
The centraliser ${\mathbf{C}}:={\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is an $F$-stable Levi subgroup of $\mathbf{G}$}. Its structure is described in [14, Exmp. 3.5.14 and 3.5.15] for $\mathbf{G}$} of classical type. Note that since $\mathbf{S}$} is a Sylow $e$-torus, it has the form ${\mathbf{T}}{\mathbf{H}}$ with $\mathbf{T}$} a torus whose order polynomial only involves factors $\Phi_{d}$ with $d\leq 2e$, and $\mathbf{H}$} a semisimple group of rank less than $e$. The claim can easily be checked by inspection of the order formulas (e.g. in [27, Tab. 24.1]). For groups of exceptional type, Table 3.3 in [14] shows that $\mathbf{C}$} is itself a torus, with the required property, unless either $e=4$ and $\mathbf{G}$} has type $E_{7}$, but the claim still holds in that case, or ${\mathbf{G}}^{F}={}^{3}\!D_{4}(q)$ and $\ell=3$ (where the claim fails). ∎
@@
Assume $\mathbf{G}$} is simple and $\Phi_{e}$ divides the order polynomial of $({\mathbf{G}},F)$. Assume the order of the relative Weyl group $W_{e}$ of a Sylow $e$-torus of $G$ is divisible by $\ell$, where $\ell>2$. Then the normaliser of a Sylow $\ell$-subgroup of $W_{e}$ contains no elements of order $\ell r$ for primes $r\equiv 1\pmod{\ell}$.
@@
Let $P\leq G$ be a Sylow $\ell$-subgroup of $G$ and set $e=e_{\ell}(q)$. By [25, Thm 5.16] there is a Sylow $e$-torus $\mathbf{S}$} of $({\mathbf{G}},F)$ such that the normaliser ${\mathbf{N}}_{G}({\mathbf{S}})$ contains the normaliser ${\mathbf{N}}_{G}(P)$ of $P$. Now ${\mathbf{L}}:={\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is an $F$-stable Levi subgroup of $\mathbf{G}$}. An application of [25, Prop. 5.3] to ${\mathbf{L}},{\mathbf{L}}]$ shows that ${\mathbf{L}},{\mathbf{L}}]^{F}$ is an $\ell^{\prime}$-group. Now ${\mathbf{L}}/[{\mathbf{L}},{\mathbf{L}}]$ is a torus, hence abelian, so as $P$ is non-abelian by assumption, $\ell$ must divide the order of the relative Weyl group $W_{e}={\mathbf{N}}_{G}({\mathbf{S}})/{\mathbf{C}}_{G}({\mathbf{S}})$.
Let $x\in G$ be an $\ell$-element. First assume $x\in{\mathbf{S}}_{\ell}^{F}$. By Proposition 5.7, since $\ell\geq 5$, $W_{e}$ has more than one Sylow $\ell$-subgroup, so $x$ lies in two different Sylow $\ell$-subgroups of ${\mathbf{N}}_{G}({\mathbf{S}})$ and hence of $G$. Now assume $x$ does not lie in any Sylow $e$-torus, and let $\mathbf{T}$} be an $F$-stable torus of $\mathbf{G}$} containing $x$. Then the order polynomial of $\mathbf{T}$} must be divisible by a cyclotomic polynomial $\Phi_{e\ell^{i}}$ for some $i\geq 1$ and hence $|{\mathbf{C}}_{G}(x)|$ is divisible by $\Phi_{e\ell^{i}}(q)$. Let $r$ be a Zsigmondy primitive prime divisor of $\Phi_{e\ell^{i}}(q)$, which exists as $e\ell^{i}$ is divisible by a prime at least 5. On the other hand, by Proposition 5.6, the order polynomial of ${\mathbf{C}}_{\mathbf{G}}({\mathbf{S}})$ is not divisible by $\Phi_{e\ell^{i}}$, so $|{\mathbf{C}}_{G}({\mathbf{S}})|$ is prime to $r$. Thus $W_{e}$ must contain an element of order $\ell r$. Now note that $r\equiv 1\pmod{\ell}$. Thus Lemma 5.8 together with Lemma 2.1 show that $x$ cannot be picky in ${\mathbf{N}}_{G}({\mathbf{S}})$ and so neither in $G$. ∎
@@
    $G={\operatorname{Sp}}_{2n}(2)$ with $3\leq n\leq 5$;
  4. (4)
    $G={\operatorname{SO}}_{2n}^{+}(2)$ with $4\leq n\leq 5$;
  5. (5)
    $G={\operatorname{SO}}_{2n}^{-}(2)$ with $4\leq n\leq 6$; or
  6. (6)
    $G=G_{2}(2)\cong{\operatorname{SU}}_{3}(3).2$, $^{3}\!D_{4}(2)$, or $F_{4}(2)$.
@@
The proof of Theorem 5.9 goes through for the prime $\ell=3$ unless either we are in one of the exceptions of [25, Thm 5.16], in an exception of Proposition 5.7, or if $\Phi_{e3^{i}}(q)$ has no primitive prime divisor for some $i\geq 1$, where again $e=e_{3}(q)$. We discuss these in turn. The exceptions from [25, Thm 5.16] are $G={\operatorname{SL}}_{3}(\epsilon q)$ with $\epsilon q\equiv 4,7\pmod{9}$ and $G=G_{2}(q)$ with $q\equiv 2,4,5,7\pmod{9}$. For $G={\operatorname{SL}}_{3}(\epsilon q)$, by Lemma 2.4 we may consider $S={\operatorname{PSL}}_{3}(\epsilon q)$ instead. The Sylow 3-subgroups of $S$ are elementary abelian of order 9, and the centraliser of an element of order 3 has structure $((q-\epsilon)\times(q-\epsilon)/3).3$. Since the Sylow 3-normalisers have the form $3^{1+2}.Q_{8}$ for the appropriate congruences, there cannot exist any picky 3-elements for $q>4$ by Lemma 2.1.
@@
Next, the exceptions of Proposition 5.7 occur precisely for the groups ${\operatorname{SL}}_{3}(\epsilon q)$, $G_{2}(q)$ and $^{3}\!D_{4}(q)$, since these are the only groups with a relative Weyl group of a Sylow $e$-torus isomorphic to ${\mathfrak{S}}_{3}$ or $W(G_{2})$. If $G={\operatorname{SL}}_{3}(\epsilon q)$ we have $3|(q-\epsilon)$ since we assume Sylow 3-subgroups of $G$ are non-abelian, and even $9|(q-\epsilon)$ by the previous paragraph. Then, by [25, Thm 5.16] the normaliser of a Sylow 3-subgroup $P$ of $G$ lies inside the normaliser ${\mathbf{N}}_{G}({\mathbf{S}})$ of a Sylow $e$-torus $\mathbf{S}$}. If $x$ is a 3-element not lying in a conjugate of ${\mathbf{S}}_{3}^{F}$, then we may conclude as in the proof of Theorem 5.9 that $x$ is not picky. If $x\in{\mathbf{S}}_{3}^{F}$ is not regular, its centraliser involves an $A_{1}$-type group, hence is not contained in ${\mathbf{N}}_{G}({\mathbf{S}})={\mathbf{S}}^{F}.{\mathfrak{S}}_{3}$ and again $x$ is not picky by Lemma 2.1. Now assume $x\in{\mathbf{S}}_{3}^{F}$ is regular. If ${\mathbf{N}}_{G}({\mathbf{S}})$ has more than one Sylow 3-subgroup, again $x$ is not picky. If ${\mathbf{N}}_{G}({\mathbf{S}})$ has a unique Sylow 3-subgroup, that is, the Sylow 3-subgroup is normal, then the elements of order three in $W_{e}$ need to centralise the $3^{\prime}$-part of ${\mathbf{S}}^{F}$ which by inspection forces ${\mathbf{S}}^{F}$ to be a 3-group. Thus $q-\epsilon$ is a power of 3 which together with $q\equiv\epsilon\pmod{9}$ forces $q=8$, $\epsilon=-1$. In the latter case, all regular $x\in{\mathbf{S}}_{3}^{F}$ are picky by direct computation, as claimed in (1).
@@
Next, if $G={}^{3}\!D_{4}(q)$ with $q\equiv\epsilon\pmod{3}$, $\epsilon\in\{\pm 1\}$, then the normaliser of a Sylow 3-subgroup is contained in a torus normaliser of the form $N=(q^{3}-\epsilon)(q-\epsilon).W(G_{2})$ (see the discussion in the proof of [25, Thm 5.14]). Now elements of order 3 in $W(G_{2})$ act like a field automorphism on the torus of order $q^{3}-\epsilon$. In particular, as soon as $q^{2}+\epsilon q+1$ has a primitive prime divisor (necessarily distinct from 3), a Sylow 3-subgroup is not normal in $N$ and we may conclude as before. This leaves $q=2$, so $G={}^{3}\!D_{4}(2)$. Here, by explicit computation, the elements of order 9 with centraliser $A_{1}(q).(q^{3}+1)$ are picky.
Finally, if there exists no primitive prime divisor for $\Phi_{e3^{i}}(q)$ with $i\geq 1$, then $i=1$, $e=q=2$. If $x\in G$ is a picky 3-element, then it has order at most 9, as elements of order 27 have centraliser order divisible by $\Phi_{18}(2)=3^{3}19$ in contradiction to Lemma 5.8. We now discuss the various types. For $G$ of classical type, all elements of order 3 lie in a Sylow $2$-torus of $G$ and thus in at least two Sylow 3-subgroups, arguing as in the proof of Theorem 5.9. So in particular, ${\mathbf{C}}_{\mathbf{G}}(x)$ must contain a torus of rational type $\Phi_{6}$. In ${\operatorname{SL}}_{n}(q)$ any centraliser order divisible by $q^{3}+1$ is in fact divisible by $(q^{6}-1)/(q-1)$, hence by 7 when $q=2$, contradicting Lemma 5.8. In the remaining groups of classical type, the primitive 9th roots of unity can occur at most once as an eigenvalue of $x$, as otherwise the centraliser contains a subgroup ${\operatorname{SU}}_{2}(8)$ and hence elements of order 7. Let $V$ be the submodule of the natural $G$-module spanned by the eigenspaces for the eigenvalues of $x$ of order at most 3 and $H$ the induced classical group on $V$. Then $H$ has the same type as $G$, respectively type ${\operatorname{SO}}^{-\epsilon}$ if $G$ has type ${\operatorname{SO}}^{\epsilon}$ (since in that case by what we showed $x$ has exactly all six primitive 9th roots of unity as eigenvalues and the induced orthogonal group on the corresponding 6-dimensional sum of eigenspaces is of minus type). Then ${\operatorname{Sub}}_{G}(x)$ will contain the subnormaliser of $x|_{V}$ (of order 3), and by direct computation the latter is all of $H$ if $H={\operatorname{SU}}_{3}(2),{\operatorname{Sp}}_{6}(2),{\operatorname{SO}}_{8}^{+}(2)$ or ${\operatorname{SO}}_{6}^{-}(2)$. As all of these have order divisible by a prime larger than 3, we see that $x$ cannot be picky in ${\operatorname{SU}}_{n}(2),{\operatorname{Sp}}_{2n}(2),{\operatorname{SO}}_{2n}^{+}(2),{\operatorname{SO}}_{2n}^{-}(2)$ when $n\geq 9,6,6,7$ respectively. By direct computation, there do exist picky elements of order 9 in all remaining cases.
Similarly, the claim for $G_{2}(2),{}^{3}\!D_{4}(2),F_{4}(2)$ and $E_{6}(2)$ follows by explicit computation in GAP. The maximal subgroup $H=Fi_{22}$ of $S={}^{2}\!E_{6}(2)$ contains a Sylow 3-subgroup of $S$. By direct computation, $H$ possesses one class of picky 3-elements, but this is fused in $S$ with one of the non-picky classes of $H$, so $S$ has no picky 3-element. The group $G=E_{7}(2)$ has six classes of elements of order 9, with centralisers
$\Phi_{2}^{2}\Phi_{6}.{}^{2}\!A_{2}(2)A_{1}(2),\ \Phi_{2}.{}^{2}\!A_{2}(8),\ \Phi_{2}\Phi_{6}.{}^{2}\!A_{3}(2)A_{1}(2),$
$\Phi_{2}\Phi_{6}.{}^{2}\!A_{4}(2),\ \Phi_{2}\Phi_{6}.{}^{3}\!D_{4}(2),\ \Phi_{2}\Phi_{6}.A_{1}(8)A_{1}(2)$
(private communication of Frank Lübeck), while all elements of order 3 lie in a Sylow 2-torus by [21]. Now the normaliser of a Sylow 3-subgroup of $G$ is contained in the normaliser of a Sylow 2-torus and thus has the form $3^{7}.N$ where $N$ is the normaliser of a Sylow 3-subgroup of the Weyl group $W(E_{7})$, whose 2-part is $2^{3}$. Since all of the above centralisers have order divisible by at least $2^{4}$, $G$ cannot have picky 3-elements by Lemma 2.1. Finally, $G=E_{8}(2)$ has four classes of elements of order 9, with centralisers
$\Phi_{2}.{}^{2}\!A_{2}(8)A_{1}(2),\ \Phi_{2}\Phi_{6}.{}^{2}\!A_{4}(2)A_{1}(2),\ \Phi_{2}\Phi_{6}.{}^{2}\!D_{5}(2),\ \Phi_{2}\Phi_{6}.{}^{3}\!D_{4}(2)A_{1}(2)$
(again computed by Frank Lübeck), and all elements of order 3 have centralisers of semisimple rank at least 7 by [21]. Since the 2-part of the normaliser of a Sylow 3-subgroup of $G$ is just $2^{4}$, we may argue as before to see that $G$ has no picky 3-elements. ∎
The case $\ell=2$ is considerably more tricky. It will by all appearance be even messier than for $\ell=3$, involving Fermat and Mersenne primes, for example. We will consider it in forthcoming work with M. Schaeffer Fry.
5.3. Subnormalisers in Suzuki and Ree groups
@@
We now assume that $\mathbf{G}$} is of type $B_{2}$, $G_{2}$ or $F_{4}$ in characteristic $p=2,3$, or 2, respectively, and $F$ is a Steinberg endomorphism of $\mathbf{G}$} such that $G={\mathbf{G}}^{F}$ is a Suzuki or Ree group. As before, $\ell$ denotes a prime distinct from the defining characteristic $p$ of $\mathbf{G}$}. We obtain the analogue of Theorems 5.3 and 5.9:
Theorem 5.11.
@@
In the above setting, let $x\in G$ be an $\ell$-element. Then $x$ is picky if and only if $\ell>3$ and $x$ is regular, while ${\operatorname{Sub}}_{G}(x)=G$ otherwise.
@@
First assume that $\ell>3$ and $x$ is regular. Then $\ell$ does not divide the order of the Weyl group of $\mathbf{G}$} and hence the Sylow $\ell$-subgroups of $G$ are abelian (see [2, Cor. 3.13, p. 259]). We may now proceed as in the proof of Theorem 5.3, replacing cyclotomic polynomials over $\mathbb{Q}$} by suitable cyclotomic polynomials over ${\mathbb{Q}}(\sqrt{p})$, with corresponding Sylow tori, for which the Sylow theorems continue to hold, see [14, 3.5.3, 3.5.4], to conclude that $x$ is picky.
We now discuss the remaining cases. All semisimple elements $x\neq 1$ of $^{2}\!B_{2}(q^{2})$ are regular and have order prime to 6, so there is nothing to prove in this case. In $G={}^{2}G_{2}(q^{2})$ the only $\ell$-elements $x\neq 1$ that are either non-regular or have $\ell\leq 3$ are the involutions, with ${\mathbf{C}}_{G}(x)=\langle x\rangle\times{\operatorname{PSL}}_{2}(q^{2})$. Such $x$ are also contained in a Sylow 2-normaliser, of structure $2^{3}.7.3$, not lying in the maximal subgroup ${\mathbf{C}}_{G}(x)$. Thus ${\operatorname{Sub}}_{G}(x)=G$ as claimed.
Let $G={}^{2}\!F_{4}(q^{2})$. The conjugacy classes of semisimple elements are given in [33, Tab. IV]. First assume $\ell>3$. The non-regular representatives $x\neq 1$ are $t_{i}$ with $i\in\{1,2,4,5,7,8\}$, where $t_{1},t_{2},t_{5}$ and $t_{8}$ only occur if $q^{2}\geq 8$. The non-cyclic Sylow $\Phi$-tori are $T(j)$ for $j\in\{1,6,7,8\}$ in [33, (3.1)], with relative Weyl groups of order $16,96,96,48$ respectively. Comparison with the list of maximal subgroups of $G$ in [23, Main Thm] shows that none of them can contain both ${\mathbf{C}}_{G}(x)$ and ${\mathbf{N}}_{G}({\mathbf{T}})$ for $\mathbf{T}$} a Sylow $\Phi$-torus of $\mathbf{G}$} containing $x$.
So finally assume $\ell=3$. Any 3-element $x\in G$ is contained in a Sylow $\Phi_{4}$-torus $\mathbf{T}$} of $\mathbf{G}$}, where ${\mathbf{N}}_{G}({\mathbf{T}})=(q^{2}+1)^{2}.G_{12}$ with the primitive complex reflection group $G_{12}$ of order 48. As ${\mathbf{N}}_{G}({\mathbf{T}})$ is a maximal subgroup of $G$ by [23, Main Thm], either ${\operatorname{Sub}}_{G}(x)={\mathbf{N}}_{G}({\mathbf{T}})$ or ${\operatorname{Sub}}_{G}(x)=G$. The structure of a Sylow 3-normaliser $N$ is discussed in [25, Thm 8.4]. If $q^{2}\equiv 8\pmod{9}$ then $N$ is contained in the torus normaliser $(q^{2}+1)^{2}.G_{12}$. Now $G_{12}$ has non-normal Sylow 3-subgroups and hence arguing as in the proof of Theorem 5.9 there cannot exist picky 3-elements, whence ${\operatorname{Sub}}_{G}(x)=G$ by what we said above. On the other hand, if $q^{2}\equiv 2,5\pmod{9}$ a Sylow 3-normaliser is isomorphic to ${\operatorname{SU}}_{3}(2).2$, not contained in any conjugate of ${\mathbf{N}}_{G}({\mathbf{T}})$ by [25, Thm 8.4(b)], so again we have ${\operatorname{Sub}}_{G}(x)=G$. ∎
By explicit computation, the subnormaliser of an element of order 3 in the Tits group $^{2}\!F_{4}(2)^{\prime}$ is a maximal subgroup ${\operatorname{PSL}}_{3}(3).2$ (not invariant under the outer automorphism of order 2).